# Correlated Systematic Uncertainties and Errors-on-Errors in Measurement Combinations with an Application to the $7$-$8$ TeV ATLAS-CMS Top Quark Mass Combination

### v3

@@ Abstract
The Gamma Variance Model (GVM) is a statistical model that incorporates uncertainties in the assignment of systematic errors (informally called errors-on-errors). The model is of particular use in analyses that combine the results of several measurements. In the past, combinations have been carried out using two alternative approaches: the Best Linear Unbiased Estimator (BLUE) method or what we will call the nuisance-parameter method. In this paper we derive useful relations that facilitate application of the model in practice. Specifically, we obtain a general relation between the BLUE and nuisance-parameter methods when the correlations induced by systematic uncertainties are non-trivial (i.e., not ±$\pm 1$ or 00), and we then generalise the nuisance-parameter approach to include errors-on-errors. We then present analytical formulas for estimating central values, confidence intervals, and goodness-of-fit when errors-on-errors are incorporated into the statistical model. To illustrate the properties of the GVM, we apply it to the 77778888 TeV ATLAS-CMS top quark mass combination. We also explore a hypothetical scenario by artificially adding a fictitious measurement as an outlier to the combination, illustrating a key feature of the GVM — its sensitivity to the internal consistency of the input data — which could become relevant for future combinations.
@@
Keywords: Gamma Variance Model, errors-on-errors, combining measurements, correlated systematics, nuisance parameters, Bartlett correction, ATLAS-CMS top mass combination
@@ 1 Introduction
In Particle Physics analyses such as those at the Large Hadron Collider (LHC), many results are already dominated by systematic uncertainties or will be shortly. In this context, uncertainties in the assigned values of systematic errors themselves – informally called errors-on-errors – become increasingly important. Often, systematic uncertainties are estimated using real datasets of control measurements or Monte Carlo simulations. However, this is not always the case. Theoretical uncertainties often require ad hoc procedures for estimation, which can carry an intrinsic degree of uncertainty. Additionally, “two-point systematics”, where uncertainties are estimated from the difference in results from two methods, provide only a limited picture of the actual uncertainty. The Gamma Variance Model (GVM) bib:Cowan2019 ; bib:Canonero2023 provides a statistical framework to handle errors-on-errors.
@@
In this paper, we focus on the application of the GVM to combinations, and we derive a number of new results that facilitate use of the model in a non-trivial analysis. We first show how a likelihood with nuisance parameters can be constructed when the correlations induced by systematic uncertainties are not simply ±$\pm 1$ or 00, and how this approach relates to the Best Linear Unbiased Estimator (BLUE) method bib:Lyons ; bib:Valassi2003 ; bib:Nisius2020 ; bib:Nisius2014 . Next, we demonstrate how this model can be generalized to include errors-on-errors and derive analytical formulas for its application under these conditions. This framework is then applied to the $7$-$8$ TeV ATLAS-CMS top quark mass combination bib:ATLAS/CMS . Specifically, we investigate the impact on the combination’s central value and confidence interval when the largest sources of systematic uncertainties are treated as uncertain. We present the results by considering the largest systematic uncertainties as uncertain, one at a time, and we vary their associated error-on-error parameters. This approach serves as a general method to identify the systematic uncertainties to which a combination is sensitive when errors-on-errors are considered.
@@
Finally, we explore how the inclusion of a hypothetical measurement that is in tension with the rest of the data affects the combination’s sensitivity to uncertain systematics. This exercise can be relevant for future combinations, as some existing top mass measurements deviate from the result of the ATLAS-CMS combined result bib:Tevatron2014 ; bib:ATLAS2023 . This example illustrates the behavior of the GVM in scenarios where the inputs of a combination are not entirely internally compatible. Unlike a standard combination, incorporating errors-on-errors makes the central value of the combination, here taken as the maximum likelihood estimator for the parameter of interest, more robust to discrepancies between input data values. Specifically, the central value becomes less sensitive to outliers. At the same time, the presence of discrepancies results in an increase in the confidence interval for the parameter of interest. This increase reflects the uncertainty stemming from the fact that the measurements in the dataset are not entirely compatible.
@@
In Sec. 2, we present an overview of how a combination can be performed using nuisance parameters under non-trivial correlation assumptions, and how this method relates to the BLUE approach. Section 3 illustrates how to extend combinations to take errors-on-errors into account. In Sec. 4, we provide analytical formulas for profiling over nuisance parameters, confidence intervals, and goodness-of-fit when errors-on-errors are incorporated into the statistical model. In Sec. 5, we apply the GVM to the 7-8 TeV ATLAS-CMS top mass combination and explore the impact of introducing an outlier to the combination. A summary and conclusions are presented in Sec. 6.
@@ 2 Equivalence between nuisance-parameter and BLUE methods in combinations with non-trivial correlations
It is usually convenient to redefine the factors $\Gamma_{i}^{s}$, the NPs $\theta_{s}$ and the auxiliary measurements $u_{s}$ such that $u_{s}$ is Gaussian distributed about $\theta_{s}$ with a variance of unity, i.e., $\sigma_{u_{s}}^{2}=1$. In the real experiment, the best estimates of the $u_{s}$ are zero. The idea here is that if there are known biases in the measurements, these are already subtracted. The remaining systematic uncertainty lies in the fact that this bias removal might be imprecise, due to the limited accuracy with which we can identify and quantify the biases. Nevertheless, here we will retain the complete notation in our discussion for greater generality. Auxiliary measurements must be explicitly included to perform Monte Carlo (MC) simulations, while the systematic variances $\sigma^{2}_{u_{s}}$ are needed to generalise the model to include errors-on-errors.
@@ 2.1 Systematics with non-trivial correlations
The BLUE method has been used, e.g., in the top-mass combination bib:ATLAS/CMS that we consider in Sec. 5. To include errors-on-errors in it, we need to first relate the BLUE likelihood of the published analysis, which contains non-trivial correlations, to an equivalent formulation in terms of nuisance parameters. It is conceptually more complicated to extend the log-likelihood containing NPs of Eq. (3) so that it gives non-trivial correlations. To accomplish this, one needs to define a distinct NP $\theta^{i}_{s}$ associated with the systematic effect $s$ for each measurement $y_{i}$, and introduce correlations between the corresponding auxiliary measurements such that
@@
Sometimes, given a systematic effect $s$, one does not need to define an individual NP $\theta^{i}_{s}$ for each measurement $i$, as many induced correlations will still be ±$\pm 1$ or 00. For example, one may have two groups of measurements A and B, and a systematic effect s which is fully correlated inside each group of measurements, but induces a $0.5$ correlation in between the two groups. In this case one would only need to introduce two NPs for the systematic effect $s$, $\theta^{A}_{s}$ and $\theta^{B}_{s}$.
Note that here we are not introducing correlations between estimates of NPs associated with different systematics, as is done in the model proposed in bib:Kieseler2017 . That model addresses fits with simultaneously constrained uncertainties, where estimates of POIs and NPs are correlated with each other. Here, instead, for each systematic $s$, we introduce a separate NP $\theta^{i}_{s}$ and auxiliary measurement $u_{s}^{i}$ for each measurement $i$, and we correlate only the auxiliary measurements for a given systematic source $s$. We do this to match the BLUE result obtained with the systematic covariance matrix defined in Eq. (8).
@@ 3 Extending combinations to account for errors-on-errors
In this section we show how to extend a combined analysis of measurements to include errors-on-errors. We first treat in Sec. 3.1 the case where the systematic uncertainties relate to those of the corresponding BLUE analysis containing only “trivial” correlations, i.e., the correlation coefficients $\rho_{ij}^{(s)}$ in the systematic covariance matrix (8) are either 00 or ±$\pm 1$. Then in Sec. 3.2 we generalise the model to cover non-trivial correlations.
@@ 3.1 The GVM with trivial correlations
The likelihoods of Eqs. (3) and (10) can be extended to account for the presence of uncertainties in the assignment of systematic errors, informally denoted as errors-on-errors. This is achieved using the Gamma Variance Model (GVM) bib:Cowan2019 ; bib:Canonero2023 . In this model the variances $\sigma^{2}_{u_{s}}$ of the control measurements $u_{s}$ are treated as adjustable parameters, and each $\sigma^{2}_{u_{s}}$ is constrained by its best estimate $v_{s}=s_{s}^{2}$. The estimates $\boldsymbol{v}$ are assumed to follow independent gamma distributions, which are suitable for positive-valued random variables. Detailed methodology and motivation for this approach can be found in bib:Cowan2019 . The resulting log-likelihood of this model is:
@@
This likelihood can be maximized in closed-form with respect to $\sigma^{2}_{u_{s}}$ for fixed $\mu$ and $\boldsymbol{\theta}$ using
@@
The parameter $\varepsilon_{s}$ quantifies the relative uncertainty associated with the systematic error $\sigma_{u_{s}}$ and is commonly referred to as the error-on-error parameter. For instance, if $\varepsilon_{s}$ equals $0.3$, this indicates that $s_{s}=\sqrt{v_{s}}$ has a 30% relative uncertainty as an estimate of $\sigma_{u_{s}}$. In most cases $\varepsilon_{s}$ parameters must be supplied by the analyst using whatever information is available on the uncertainties of $\sigma_{u_{s}}$. In the absence of the required expert knowledge to inform such a choice, results of the GVM may be presented as a function of assumed values for $\varepsilon_{s}$. The model has also been discussed in bib:Reader2024 in the context of Parton Distribution Functions (PDF) fitting.
@@ 3.2 The GVM with non-trivial correlations
The likelihood of Eq.(10) is designed to admit non-trivial correlations, i.e., not only 00 or ±$\pm 1$. Generalising it to include errors-on-errors requires extending the GVM beyond what was originally presented in bib:Cowan2019 . The method to achieve this is nevertheless similar to what has already been discussed: the variance $\sigma_{u_{s}}^{2}$ in Eq. (10) is treated as an adjustable parameter of the model, and $v_{s}$, its best estimate, is considered a gamma-distributed random variable. The resulting generalised log-likelihood is
@@
This likelihood can be maximized with respect to $\sigma^{2}_{u_{s}}$ using
@@
Often, in a manner similar to that of Sec. 2, it is useful to redefine $\Gamma_{s}$, $\theta_{s}$ and $u_{s}$ such that if there are no errors-on-errors ($\sigma_{u_{s}}^{2}$ exactly known), then $u_{s}$ is Gaussian distributed with a mean $\theta_{s}$ and unit variance. In this way the best estimates of the variances are all $v_{s}=1$. This does not alter the approach used to include errors-on-errors, as it is merely a redefinition of parameters.
In practical applications, the most feasible approach to account for errors-on-errors involves using NPs exclusively for systematic sources where $\varepsilon_{s}\neq 0$, and profiling overall NPs introduced as standard quadratic constraints. This leads to a mixed likelihood with a quadratic part constructed using the BLUE approach, and with NPs introduced only for systematic sources with uncertain errors. This can be achieved by replacing $V$ with a BLUE covariance matrix $W$, as defined in Eq. 5, but summing only over the non-uncertain systematic errors while keeping in the likelihood only those NPs associated with uncertain systematics. This can be done in a similar manner to the proof of the equivalence between the NPs approach and the BLUE approach, as shown in App. A.
Notice that when the off-diagonal correlation coefficients $\rho^{(s)}_{ij}$ of Eq. (9) are set to zero for some systematic effect $s$, the logarithmic term for that systematic effect reduces to
$\left(N+\frac{1}{2\varepsilon^{2}_{s}}\right)\log\left[1+2\varepsilon_{s}^{2}\sum_{i=1}^{N}\frac{(u^{i}_{s}-\theta^{i}_{s})^{2}}{v_{s}}\right]\,.$ (17)
This result stems from the assumption that all $u_{s}^{i}$ for a fixed $s$ share the same variance $\sigma_{u_{s}}^{2}$, with a single estimated value $v_{s}$. In practice, this means assuming that if one systematic uncertainty is overestimated or underestimated by a certain factor, this factor will be the same for all the measurements affected by that systematic. Therefore, this model is suitable for modeling systematic uncertainties of the same type or those estimated using similar procedures.
If one wants to consider both the auxiliary measurements as uncorrelated and the estimates of the systematic uncertainties as independent, one would define the logarithmic terms as
$\sum_{i=1}^{N}\left(1+\frac{1}{2\varepsilon^{2}_{s}}\right)\log\left[1+2\varepsilon_{s}^{2}\frac{(u^{i}_{s}-\theta^{i}_{s})^{2}}{v^{i}_{s}}\right]\,.$ (18)
In this case, the original GVM, as defined in Eq. (13), is recovered, with the systematic uncertainty $s$ treated as independent for each measurement $i$.

### v4

@@ Abstract
The Gamma Variance Model is a statistical model that incorporates uncertainties in the assignment of systematic errors (informally called errors-on-errors). The model is of particular use in analyses that combine the results of several measurements. In the past, combinations have been carried out using two alternative approaches: the Best Linear Unbiased Estimator (BLUE) method or what we will call the nuisance-parameter method. In this paper, we obtain a general relation between the BLUE and nuisance-parameter methods when the correlations induced by systematic uncertainties are non-trivial (i.e., not ±$\pm 1$ or 00), and we then generalise the nuisance-parameter approach to include errors-on-errors. We then present analytical formulas for estimating central values, confidence intervals, and goodness-of-fit when errors-on-errors are incorporated into the statistical model. To illustrate the properties of the Gamma Variance Model, we apply it to the 77778888 TeV ATLAS-CMS top quark mass combination. We also explore a hypothetical scenario by artificially adding a fictitious measurement as an outlier to the combination, illustrating a key feature of the Gamma Variance Model — its sensitivity to the internal consistency of the input data — which could become relevant for future combinations.
@@
Keywords: Gamma Variance Model, errors-on-errors, combining measurements, correlated systematics, nuisance parameters, Bartlett correction, ATLAS-CMS top-mass combination
@@ 1 Introduction
In Particle Physics analyses such as those at the Large Hadron Collider (LHC), many results are already dominated by systematic uncertainties or will be shortly. In this context, uncertainties in the assigned values of systematic errors themselves – informally called errors-on-errors – become increasingly important. Often, systematic uncertainties are estimated using real datasets of control measurements or Monte Carlo simulations. In such cases, the limited size of these datasets introduces an uncertainty in the assigned value of the systematic error, which has a statistical nature. However, this is not the only source of uncertainty in its assigned value. For some systematic uncertainties, the definition of a “$1$-$\sigma$” error can itself be ambiguous. Theoretical uncertainties, for example, often rely on ad hoc estimation procedures that inherently carry an additional degree of uncertainty. Furthermore, “two-point systematics”, where uncertainties are estimated from the difference in results from two methods, provide only a limited picture of the actual uncertainty. The Gamma Variance Model (GVM) bib:Cowan2019 ; bib:Canonero2023 offers a statistical framework to handle errors-on-errors of both types.
@@
In this paper, we focus on the application of the GVM to combinations, and we derive a number of new results that facilitate the use of the model in a non-trivial analysis. We first show how a likelihood with nuisance parameters can be constructed when the correlations induced by systematic uncertainties are not simply ±$\pm 1$ or 00, and how this approach relates to the Best Linear Unbiased Estimator (BLUE) method bib:Lyons ; bib:Valassi2003 ; bib:Nisius2020 ; bib:Nisius2014 . Next, we demonstrate how this model can be generalized to include errors-on-errors and derive analytical formulas for its application under these conditions. This framework is then applied to the $7$-$8$ TeV ATLAS-CMS top quark mass combination bib:ATLAS/CMS . Specifically, we investigate the impact on the combination’s central value and confidence interval when the largest sources of systematic uncertainties are treated as uncertain. We present the results by considering the largest systematic uncertainties as uncertain, one at a time, and we vary their associated error-on-error parameters. This approach serves as a general method to identify the systematic uncertainties to which a combination is sensitive when errors-on-errors are considered.
@@
Finally, we explore how the inclusion of a hypothetical measurement that is in tension with the rest of the data affects the combination’s sensitivity to uncertain systematics. This exercise can be relevant for future combinations, as some existing top-mass measurements deviate from the result of the ATLAS-CMS combined result bib:Tevatron2014 ; bib:ATLAS2023 . This example illustrates the behavior of the GVM in scenarios where the inputs of a combination are not entirely internally compatible. Unlike a standard combination, incorporating errors-on-errors makes the central value of the combination, here taken as the maximum likelihood estimator for the parameter of interest, more robust to discrepancies between input data values. Specifically, the central value becomes less sensitive to outliers. At the same time, the presence of discrepancies results in an increase in the confidence interval for the parameter of interest. This increase reflects the uncertainty stemming from the fact that the measurements in the dataset are not entirely compatible.
@@
In Sec. 2, we present an overview of how a combination can be performed using nuisance parameters under non-trivial correlation assumptions, and how this method relates to the BLUE approach. Section 3 illustrates how to extend combinations to take errors-on-errors into account. In Sec. 4, we provide analytical formulas for profiling over nuisance parameters, and computing confidence intervals and goodness-of-fit when errors-on-errors are incorporated into the statistical model. In Sec. 5, we apply the GVM to the 7-8 TeV ATLAS-CMS top-mass combination and explore the impact of introducing an outlier to the combination. A summary and conclusions are presented in Sec. 6.
@@ 2 Equivalence between nuisance-parameter and BLUE methods in combinations with non-trivial correlations
It is usually convenient to redefine the factors $\Gamma_{i}^{s}$, the NPs $\theta_{s}$ and the auxiliary measurements $u_{s}$ such that $u_{s}$ is Gaussian distributed about $\theta_{s}$ with a variance of unity, i.e., $\sigma_{u_{s}}^{2}=1$. With this choice, the value of $\Gamma_{i}^{s}$ corresponds to the systematic error induced by source $s$ in the measurement $y_{i}$. Furthermore, in the real experiment, the best estimates of the $u_{s}$ are zero. The idea here is that if there are known biases in the measurements, these are already subtracted. The remaining systematic uncertainty lies in the fact that this bias removal might be imprecise, due to the limited accuracy with which we can identify and quantify the biases. Nevertheless, here we will retain the complete notation in our discussion for greater generality. Auxiliary measurements must be explicitly included to perform Monte Carlo (MC) simulations, while the systematic variances $\sigma^{2}_{u_{s}}$ are needed to generalise the model to include errors-on-errors.
@@ 2.1 Systematics with non-trivial correlations
It is conceptually more complicated to extend the log-likelihood containing NPs of Eq. (3) so that it gives non-trivial correlations. To accomplish this, one needs to define a distinct NP $\theta^{i}_{s}$ associated with the systematic effect $s$ for each measurement $y_{i}$, and introduce correlations between the corresponding auxiliary measurements such that
@@
2.2 Comparison with the Convino framework
Correlations between auxiliary measurements in combinations are implemented within the software package proposed in bib:Kieseler2017 called Convino. In this framework, auxiliary measurements associated with different sources of systematic uncertainties can be treated as correlated. This approach leads to a systematic constraint term, expressed in our notation as
$-\frac{1}{2}\sum_{s=1}^{M}(u_{s}-\theta_{s})C_{sp}^{-1}(u_{p}-\theta_{p})\,,$ (11)
where $C$ is the covariance matrix representing the correlations among auxiliary measurements associated with different systematic effects. This matrix is crucial for handling fits with simultaneously constrained uncertainties, where the NP estimates are correlated.
In contrast, our approach focuses on extending the NP basis for each individual systematic source $s$ to induce the desired correlation structure in the measurements $\boldsymbol{y}$. However, we do not introduce correlations between different sources of systematic uncertainties, i.e., auxiliary measurements with different $s$ indices remain independent. Specifically, for each systematic $s$, we introduce a separate NP $\theta_{s}^{i}$ and auxiliary measurement $u_{s}^{i}$ for each measurement $i$. This leads to correlations only between auxiliary measurements that share the same $s$ index but have different $i$ indices, and results in a systematic constraint term given by
$-\frac{1}{2}\sum_{s=1}^{M}\sum_{i,j=1}^{N}(u^{i}_{s}-\theta^{i}_{s})\frac{1}{\sigma_{u_{s}}^{2}}\left(\rho^{(s)}\right)_{ij}^{-1}(u^{j}_{s}-\theta^{j}_{s})\,.$ (12)
This extension is necessary to achieve complete equivalence between the BLUE method and the nuisance-parameter approach under any correlation assumption.
After introducing new NPs $\theta_{s}^{i}$ for a given source $s$, the Convino framework offers a natural, and more general, extension by allowing correlations between different systematic effects. More specifically, it allows for correlations among auxiliary measurements that have distinct $s$ indices. Nevertheless, by mapping $s$ and $i$ onto a single index using a mapping function $f(s,i)$, the Convino framework can be applied in its current form to incorporate the NP approach discussed here. However, the formalism we use clearly highlights the connection between different independent systematic sources and the correlation structure they induce in the measurements $\boldsymbol{y}$, as defined within the framework of the BLUE method.
@@ 3 Extending combinations to account for errors-on-errors
In this section, we demonstrate how to extend a combination to account for the possibility that a systematic uncertainty may itself be uncertain. These uncertainties, referred to as errors-on-errors, can arise from ambiguities in the definition of a $1-\sigma$ error, as is often the case with theoretical uncertainties or two-point systematics. Additionally, they can result from the estimation of a systematic effect based on a limited sample of MC events or real data. Accounting for such uncertainties can have a non-negligible, or even significant, impact on the combination results, as demonstrated in the example provided in Sec. 5.
@@ 3.1 The GVM with trivial correlations
The likelihood of Eq. (3) can be extended to account for the presence of uncertainties in the assignment of systematic errors, informally denoted as errors-on-errors. This is achieved using the Gamma Variance Model (GVM) bib:Cowan2019 ; bib:Canonero2023 . This model allows the systematic uncertainty $\sigma_{u_{s}}$ to be treated as uncertain. Specifically, its square, the variance $\sigma_{u_{s}}^{2}$, is treated as an adjustable parameter of the model, effectively introducing it as an NP. For mathematical convenience, we use the variance $\sigma^{2}_{u_{s}}$ as the model’s NP rather than the standard deviation $\sigma_{u_{s}}$. In the GVM, the value that would typically be assigned to $\sigma^{2}_{u_{s}}$ in a standard combination is treated as an estimate of this parameter and is denoted as $v_{s}$.
One should keep in mind that the choice of independent gamma distributions for the estimates $\boldsymbol{v}$ is not expected to produce a perfect model, rather an approximation that represents a potentially significant improvement over the assumption of exactly specified variances. Other distributions, provided they are defined for positive variables, could also be considered. Some studies carried out in connection with bib:Cowan2019 of the log-normal distribution or a gamma distribution for the estimated standard deviations led to qualitatively similar results, but were not pursued further when the following reasons to use gamma-distributed $\boldsymbol{v}$ became apparent.
First, suppose that $u_{s}$ itself is found as an average of some number of other Gaussian-distributed values $u_{s,i}$, $i=1,...,n$, and the estimate of $\sigma^{2}_{u_{s}}$ is found as the sample variance of these $n$ values. In this case, the estimated variance follows a gamma distribution. In general, the $u_{s}$ might not be found in this way and the gamma distribution therefore is not expected to provide an exact model. The scenario of multiple Gaussian measurements leading to $u_{s}$ is nevertheless plausible and provides an important point of contact with our model assumptions.
Second, the gamma distribution provides significant mathematical simplifications, as will be demonstrated in this section. Detailed methodology and motivation for this approach can be found in bib:Cowan2019 .
The log-likelihood of GVM is defined as
@@
This model extends Eq. (3) by adding, in the last line, terms associated with treating $\sigma^{2}_{u_{s}}$ as an adjustable parameter of the model and its estimate $v_{s}$ as gamma distributed. The parameter $\varepsilon_{s}$ quantifies the relative uncertainty associated with the systematic error $\sigma_{u_{s}}$ and is commonly referred to as the error-on-error parameter. For instance, if $\varepsilon_{s}$ equals $0.3$, this indicates that $s_{s}=\sqrt{v_{s}}$ has a 30% relative uncertainty as an estimate of $\sigma_{u_{s}}$.
In principle, the $\varepsilon_{s}$ parameters should be provided by the analyst based on any available information regarding the uncertainties in $\sigma_{u_{s}}$. It can represent the statistical uncertainty of $\sigma_{u_{s}}$ arising from the limited size of the auxiliary measurement dataset or be assigned using more ad hoc arguments for uncertainties such as theoretical errors. In the absence of the required expert knowledge to inform such a choice, results of the GVM may be presented as a function of assumed values for $\varepsilon_{s}$.
One of the appealing mathematical properties of this model is that the log-likelihood can be maximized in closed form with respect to $\sigma^{2}_{u_{s}}$, for fixed $\mu$ and $\boldsymbol{\theta}$, using
@@
In this discussion, we denote $\sigma_{u_{s}}$ the standard deviation of $u_{s}$ as the systematic error. However, in a manner similar to that of Sec. 2, it is often convenient to redefine $\Gamma^{s}_{i}$, $\theta_{s}$, and $u_{s}$ so that, in the absence of errors-on-errors (i.e., when $\sigma_{u_{s}}$ is precisely known), the variable $u_{s}$ is Gaussian-distributed with mean $\theta_{s}$ and unit variance. Consequently, the best estimates of all variances become $v_{s}=1$. Under this choice, the actual estimate of the systematic error induced by source $s$ in measurement $y_{i}$ is expressed as $\Gamma_{i}^{s}$. Thus, one simply needs to fix the $\Gamma_{i}^{s}$ factors to the systematic uncertainty values. This redefinition does not affect the method of introducing errors-on-errors, as the model now accounts for potential fluctuations in $v_{s}$, which rescale $\Gamma_{i}^{s}$ whenever the systematic error is over- or underestimated.
The inclusion of errors-on-errors replaces the standard quadratic constraints of Eq. (3), typically used to model systematic effects, with the logarithmic constraints of Eq. (15). These logarithmic terms are quadratic and identical to those in Eq. (3) when $\varepsilon_{s}$ approaches zero. However, for larger values of $\varepsilon_{s}$, these log-terms penalize the log-likelihood less for values of $\theta_{s}$ that deviate from $u_{s}$, thus changing the way the model handles incompatible data. An example of this property will be provided in Sec. (5.3). An application of this model has also been presented in bib:Reader2024 in the context of Parton Distribution Function (PDF) fitting.
@@ 3.2 The GVM with non-trivial correlations
The likelihood of Eq.(10) is designed to admit non-trivial correlations, i.e., not only 00 or ±$\pm 1$. This is achieved by extending the NP basis for each systematic source $s$, introducing a distinct NP $\theta_{s}^{i}$ and auxiliary measurement $u_{s}^{i}$ for each measurement $i$, with correlations applied only among auxiliary measurements sharing the same $s$ index but different $i$ indices.
Generalising it to include errors-on-errors requires extending the GVM beyond what was originally presented in bib:Cowan2019 . The method to achieve this remains similar to the previously discussed approach: the variances $\sigma_{u_{s}}^{2}$ in Eq. (10) are treated as adjustable parameters within the model, meaning their exact values are now considered unknown. Meanwhile, their best estimates, $v_{s}$, are modeled as independently gamma-distributed random variables and incorporated into the model. The resulting log-likelihood is
@@
Here, the terms in the last line arise from treating $\sigma^{2}_{u_{s}}$ as an adjustable parameter of the model and $v_{s}$, its estimate, as gamma-distributed. This log-likelihood can be maximized with respect to $\sigma^{2}_{u_{s}}$ using
@@
In this model, one would typically redefine the parameters in such a way that the estimate $v_{s}$ is fixed to one. For a given source $s$, any overestimation or underestimation of the systematic uncertainties induced in the measurements $\boldsymbol{y}$ occurs uniformly by the same factor.
3.3 Relationship between the two GVM versions
In this section, we introduced two methods for incorporating errors-on-errors. To clarify their differences, consider a simplified example with two systematic effects, $A$ and $B$. In the first approach, the auxiliary measurements $u_{A}$ and $u_{B}$ have variances $\sigma^{2}_{u_{A}}$ and $\sigma^{2}_{u_{B}}$, with their best estimates $v_{A}$ and $v_{B}$ treated as independent random variables. This means the statistical fluctuations of $v_{A}$ and $v_{B}$ are uncorrelated, allowing scenarios where $v_{A}$ underestimates $\sigma^{2}_{u_{A}}$ while $v_{B}$ overestimates $\sigma^{2}_{u_{B}}$, or vice versa, or where both over- or underestimate their respective variances. These assumptions result in the following logarithmic constraints:
$\begin{split}&-\frac{1}{2}\left(1+\frac{1}{2\varepsilon_{A}^{2}}\right)\log\left[1+2\varepsilon_{A}^{2}\frac{(u_{A}-\theta_{A})^{2}}{v_{A}}\right]\\ &-\frac{1}{2}\left(1+\frac{1}{2\varepsilon_{B}^{2}}\right)\log\left[1+2\varepsilon_{B}^{2}\frac{(u_{B}-\theta_{B})^{2}}{v_{B}}\right]\,,\end{split}$ (19)
where $\varepsilon_{A}$ and $\varepsilon_{B}$ are the errors-on-errors parameters. This construction defines the model in Eq. (15) that we use to handle trivial correlations.
In the second approach, $v_{A}$ and $v_{B}$ are assumed to be fully dependent, meaning they either both underestimate or both overestimate their respective variances by the same factor. This is modeled by setting $\sigma^{2}_{u_{A}}=\sigma^{2}\Gamma^{2}_{A}$ and $\sigma^{2}_{u_{B}}=\sigma^{2}\Gamma^{2}_{B}$, where $\sigma^{2}$ is an unknown global variance parameter, shared by both systematic effects. The best estimate of $\sigma^{2}$ is denoted by $v$, with an associated error-on-error parameter $\varepsilon$. This approach leads to the following logarithmic term:
$\small-\frac{1}{2}\left(N+\frac{1}{2\varepsilon^{2}}\right)\log\left[1+2\varepsilon^{2}\frac{(u_{A}-\theta_{A})^{2}}{v\Gamma^{2}_{A}}+2\varepsilon^{2}\frac{(u_{B}-\theta_{B})^{2}}{v\Gamma_{B}^{2}}\right]\,,$ (20)
where $N=2$, as the logarithm includes terms corresponding to two auxiliary measurements. The scaling factors $\Gamma^{2}_{A}$ and $\Gamma^{2}_{B}$ can be absorbed into the definitions of the NPs and auxiliary measurements for simplicity. Importantly, when $\varepsilon=0$, this approach becomes equivalent to the first method. Additionally, this method enables the introduction of a correlation factor between the auxiliary measurements $u_{A}$ and $u_{B}$, while also accounting for a global error-on-error affecting the scale of their variances.
The model defined in Eq. (18) to handle non-trivial correlations incorporates both assumptions. It treats the variances of auxiliary measurements from different systematic sources (indexed by $s$) as independent. In contrast, for a given source $s$, the variances of auxiliary measurements (indexed by $i$) are considered dependent, as they are associated with the same systematic effect. That is, the estimates of the variances associated with a given systematic effect $s$ are proportional to $v_{s}$.